OpenStudy (anonymous):

find the value of 256^0.16 *256^0.09

OpenStudy (anonymous):

\[256^{0.16}\times 256^{0.09}=256^{0.16+0.09}=256^{\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

\(256=2^8\). Therefore, \(256^{\frac{ 1 }{ 4 }}=\left( 2^{8} \right)^{\frac{ 1 }{ 4 }}=2^{8\times \frac{ 1 }{ 4 }}=2^2=4\)